OpenStudy (anonymous):

Please help~ ( I've been working on this problem for like a hour) triangle PQR has vertices P (4,-1),Q(-2,7), and R (9,9) b) find an equation (in standard form) of the altitude from R

OpenStudy (anonymous):



OpenStudy (anonymous):

First find slope of PQ

OpenStudy (anonymous):

-4/3

OpenStudy (anonymous):

Now, For perpendicular lines, m1 * m2 =-1

OpenStudy (anonymous):

Where m1 and m2 are the slopes of the perpendicular lines

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

idk xD

OpenStudy (anonymous):

Product of slope of perpendicular lines are -1

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

So the slope of altitude from R is 3/4

OpenStudy (anonymous):

Now, use slope-point formula

OpenStudy (anonymous):

oh I get it now

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

welcomxx

OpenStudy (anonymous):

wait but what point do I use?

OpenStudy (radar):

I would think you would now calculate the distance from point R(9,9) to the intersection of the perpendicular line and the line segment QP at the point where the x and y values of each line (the line QP and the line with the slope of 3/4 from point R.

OpenStudy (radar):

are the same.

OpenStudy (radar):

You have these two lines: y=-(4/3)x + 13/3 and y=(3/4)x + 9/4 Now find the point on each line where the x values are the same and the y values are the same.